OpenStudy (anonymous):

What is the equation of the axis of symmetry of the graph of y=ax^2+bx+c ? a. y= -b/2a b. x= -b/2a c. y= -b/2c d. x= -b/2c

OpenStudy (ranga):

The x-coordinate of the vertex of the equation: y = ax^2 + bx + c is given by x = -b/2a. The vertical line through -b/2a will be the axis of symmetry. So the equation of the vertical line is x = - b / 2a